# The non-linear dynamics of axion inflation: a detailed lattice study

### v1

@@ IV.1 Initial condition on the lattice
In order to run any simulation, we need first to initialize all the fields in the lattice. In the case of the gauge field, we know that the Bunch-Davies (BD) vacuum (23) is the solution that sets the initial condition for each helicity component, for wavelengths well inside the Hubble radius. Such condition characterizes an initial quantum vacuum state $|0\rangle$ (for which $\hat{a}_{\bf k}|0\rangle=0$), whereas solving for Eqs. (31)-(33) [or more precisely for their lattice counterparts Eqs. (44)-(45) and (47)] assumes implicitly the use of classical fields. Given the quantum expectation value $s_{A}^{2}(\tau,{\mathbf{k}})\equiv\langle 0|\hat{A}^{\lambda}(\tau,{\bf k})\hat{A}^{\lambda^{\dagger}}(\tau,{\bf k})|0\rangle\equiv|\mathcal{A}^{\lambda}(\tau,{\bf k})|^{2}$ for the chirality $\lambda$, the BD condition implies $s_{A}^{2}\longrightarrow{1\over 2k}$. This corresponds to initially vanishing occupation numbers, $n_{\bf k}=0$, however, we know that one of the gauge field chiralities ($\lambda=+$ in our convention) will grow exponentially fast during the linear regime. This leads, in turn, to exponentially growing occupation numbers, $n_{\bf k}\gg 1$, and as a result, we can actually approximate the vacuum expectation values of products of field operators, by ensemble averages of random fields, see e.g. [98]. This corresponds to the classical limit of the excited field(s).
@@
$A^{\pm}(\mathcal{N},\tilde{\bf n})$ $\simeq$ $f^{\pm}_{k}e^{ik/(aH)}\,,$ (61)
$E^{\pm}(\mathcal{N},\tilde{\bf n})$ $\simeq$ $-\frac{i}{a}kg_{k}^{\pm}e^{ik/(aH)}\,,$ (62)
where $f^{+}_{k},f^{-}_{k},g^{+}_{k},g^{-}_{k}$ are the independent amplitude realizations of a Gaussian random field with zero mean and root-mean-square ${1/\sqrt{2k}}$. We note that $\mathcal{N}$ appears only implicitly throughout $a=a(\mathcal{N})$ and $H=H(\mathcal{N})$, and we highlight that $k$ in Eqs. (61)-(62) corresponds to the modulus of the linear momentum, $k=|{\mathbf{k}}|=k_{\rm IR}|\tilde{\bf n}|$, to mimic correctly the continuum spectra.
Using Eq. (55), we finally convert the chiral fluctuations (61)-(62) into Cartesian components of the gauge and the electric field, $A_{i}(\mathcal{N},{\bf n})$ and $E_{i}(\mathcal{N},{\bf n})$, where we highlight that the angles $\theta_{\rm L},\varphi_{\rm L}$ in Eq. (55) correspond to the polar and azimuthal angles of the lattice momentum ${\bf k}_{\rm L}(\tilde{\bf n})$, the Cartesian components of which are given in Eq. (52). A more detailed discussion of the implementation of the Bunch-Davies vacuum in the lattice can be found in Appendices B and C.
@@ IV.2 Linear regime on the lattice
$\bullet$ An educated guess of the window $[k_{\rm IR},k_{\rm UV}]$ needed to capture at least the modes for which the gauge field power spectrum is significantly above the BD tail, assuming linear evolution of such modes until the end of slow-roll inflation.
@@
It can be observed that initially there is a phase where the BD tail dominates over the tachyonic excitation in $\rho_{\rm EM}$. While $\rho_{\text{EM}}$ decreases with the expansion, IR modes also get excited, and eventually the excited IR range of the spectrum comes to dominate over the UV vacuum tail. In this example, this happens around $\mathcal{N}\sim-3$. Before that, the relative difference between methods is of the order of $\sim 0.1\%$, which is caused by the randomness of the BD condition in the lattice. We have checked that increasing the accuracy of the integrator does not reduce this relative error. Once the excited part of the gauge spectrum dominates over the vacuum tail, the relative difference increases, reaching approximately $\sim 10\,\%$ depending on the random realization. This is due to the cosmic variance problem already noted for the difference between spectra in the IR modes in Fig. 1. We highlight this issue by including the $1\sigma$ deviation in shaded, and we show that the variability of the fluctuation in the IR modes are causing this effect. Extending the evolution to $\mathcal{N}=0$ (while assuming inflaton slow-roll dominance), the mean relative error decreases again down to $\sim 1\,\%$, as the peak shifts to intermediate scales in the range $[k_{\text{IR}},k_{\text{UV}}]$, and the weight of the most discrepant IR bins becomes less significant. In summary, regardless of the seed for the BD fluctuations, we correctly capture the evolution of the electromagnetic energy density in the lattice.
@@ IV.3 Switching to non-linear dynamics
For the largest couplings considered, in order to prevent the unwanted contribution of the gauge field vacuum tail to affect the dynamics, an additional step is required. Due to the special sensitivity to the UV of the non-linear regime, the lattice maximum momentum $k_{\rm UV}$ (or the comoving spatial resolution) needs to be sufficient to cover all relevant modes that will be excited in such regime. If the BD vacuum solution was imposed to all those modes, the non-classical contribution to the electromagnetic backreaction terms would be too large and would become the evolution unphysical. We tackle this issue introducing an intermediate cutoff $k_{\text{IR}}<k_{\text{BD}}<k_{\text{UV}}$, so that only within the ‘IR range’ $[k_{\rm IR},k_{\rm BD}]$ we set the BD initial condition for the modes that are expected to be excited during the linear regime. Within the ‘UV range’ $[k_{\rm BD},k_{\rm UV}]$, instead, the gauge field and electric field modes are set to zero during the linear dynamics, as they are not expected to become excited from the tachyonic instability. The gauge field modes within $[k_{\rm BD},k_{\rm UV}]$ grow only stimulated out of the non-linear dynamics, when $\mathcal{N}>\mathcal{N}_{\rm switch}$. As a side effect of this procedure we can also tolerate to choose a larger time step $d\mathcal{N}$. We have performed numerical checks in order to ensure that applying such a cutoff does not alter the subsequent non-linear evolution, see App. D for a detailed discussion.
@@
The different purple lines correspond to different integration ranges for the integrated electromagnetic contributions. The dark solid lines correspond to the contribution calculated in the range $[k_{\text{IR}},k_{\text{BD}}]$ and the light ones in $[k_{\text{IR}},k_{\text{UV}}]$. While the former is the contribution we measure in our simulations, the latter is what we would have if we set the BD solution for the whole dynamical range. The non-continuous lines correspond to the isolated contribution of the IR excitation, that is above the BD initial condition of the gauge spectra. We do this by integrating the spectra up to a time-increasing cutoff $k_{\rm cut}(\mathcal{N})$ that grows from $k_{\rm IR}$ up to $k_{\rm BD}$. The dashed lines correspond to the range $[k_{\rm IR},k_{\rm cut}(\mathcal{N})]$, whereas the dotted lines to $[k_{\rm min},k_{\rm cut}(\mathcal{N})]$, where $k_{\text{min}}$ corresponds to Hubble crossing of the CMB scales at $\mathcal{N}=-60$, and with which we capture using the 1D method the gauge excitation during the whole cosmic history.
@@
Additionally, we see that the IR excitation captured by our lattice momentum window (dashed lines) is initially very subdominant, but it eventually becomes comparable to the vacuum tail’s contribution around $\mathcal{N}=-1.5$. Furthermore, it is also necessary that by the time we reach $\mathcal{N}_{\text{switch}}$, the IR excitation on the lattice dominates over that corresponding to modes below our IR threshold, i.e. $k<k_{\text{IR}}$. We assess this by including the contribution of the excited part computed in the range $[k_{\rm min},k_{\rm cut}(\mathcal{N})]$ (dotted lines). The comparison with the dashed lines demonstrates that by the time the tachyonic excitation dominates over the UV tail, it also acquires well the level of excitation that considers the full cosmic history.
Figure 3: Comparison for $\alpha_{\Lambda}=15$ (left) and $18$ (right) of the terms related to the backreaction used to establish the time $\mathcal{N}_{\text{switch}}$. Top: the evolution of $\rho_{\rm K}$ (red), $\rho_{\rm V}$ (black) and $\rho_{\rm EM}$ (purple). Bottom: the evolution of $|\dot{\pi}_{\phi}|$ (red), $3H|\pi_{\phi}|$ (orange), and $m^{2}|\phi|$ (black) of Eq. (5), together with $\frac{\alpha_{\Lambda}}{a^{3}m_{p}}|\langle\vec{E}\cdot\vec{B}\rangle|$ (purple). The different purple lines represent different integration ranges: the dark solid purple lines are the contribution integrated for the range $[k_{\text{IR}},k_{\text{BD}}]$, the light purple the integral for $[k_{\text{IR}},k_{\text{UV}}]$, the dashed line is for only the excited IR contribution for $[k_{\rm IR},k_{\rm cut}(\mathcal{N})]$, and the dotted purple line is the contribution of the excitation considering the whole cosmic history, i.e. $[k_{\rm min},k_{\rm cut}(\mathcal{N})]$ (see text). The vertical dashed gray lines correspond to the times $\mathcal{N}_{\text{switch}}$.
@@
Finally, we note that the effect of the action of the cutoff scale $k_{\rm BD}$ is more pronounced for the largest couplings. As in the previous example, we include the electromagnetic components computed considering that we set the BD vacuum solution all across the range $[k_{\text{IR}},k_{\text{UV}}]$ using light solid purple lines. It clearly shows that without the cutoff at $k_{\rm BD}$ an incredible overestimation of the integrated contributions will occur, which would falsify completely the dynamics. We conclude that for the largest coupling the use of an intermediate cutoff is compulsory.
@@ IV.4 Homogeneous Backreaction
$\alpha_{\Lambda}$ $N\ (\geq)$ $k_{\rm{IR}}/m$ $k_{\rm{UV}}/m\ (\geq)$ $k_{\rm{BD}}/m$ $\mathcal{N}_{\rm{start}}$ $\mathcal{N}_{\rm{switch}}$
@@
Table 1: Summary of the relevant parameters of our simulations, with $\alpha_{\Lambda}$ the dimensionless coupling constant, $N$ the number of lattice sites per dimension, $k_{\rm{IR}}$ the infrared momentum, $k_{\rm{UV}}$ the ultraviolet momentum, $k_{\rm{BD}}$ the intermediate cutoff scale, $\mathcal{N}_{\rm{start}}$ the efold at which we start the simulation in the linear regime and $\mathcal{N}_{\rm{switch}}$ the efold in which we evolve the full equations of motion. For each coupling, only the cases with the largest separation of scales (or $N$) are included.
@@
${\rm Linear}:$ $\Delta\mathcal{N}_{\rm{br}}=m_{2}(\alpha_{\Lambda}-14.31)+1.44\,,$ (79)
${\rm Power-law}:$ $\Delta\mathcal{N}_{\rm{br}}=b_{2}(\alpha_{\Lambda}-14.31)^{a_{2}}+1.44\,,$ (80)
@@
[…]
@@
[…]
@@ V.2 Inflating UV scales
We observe that although there is a convergence in $\epsilon_{H}$ in the initial phase of backreaction, $-2\lesssim\mathcal{N}\lesssim 2$, trajectories start diverging during the final stages towards the end of inflation. Simulations with the worst resolution (smaller $k_{\rm UV}/k_{\rm IR}$ ratio) deviate earlier, whereas the ones with better resolution sustain convergence up to $\mathcal{N}\sim 3.5$, and then diverge afterwards.
@@ V.3 Partial restoration of chirality balance and loss of transversality
$\big{(}\pi_{\phi}\vec{B}\,\big{)}_{\vec{k}}=-i\sum_{\lambda=\pm}\lambda\int d^{3}q\,\,\dot{\phi}^{*}_{(\vec{k}\shortminus\vec{q})}\,q\,A^{\lambda}_{\vec{q}}\,\vec{\varepsilon}^{\,\lambda}_{\vec{q}}\,,$ (84)
$\big{(}\vec{\nabla}\phi\times\vec{E}\,\big{)}_{\vec{k}}=i\sum_{\sigma=\pm\,,\parallel}\int d^{3}q\,\,\phi^{*}_{(\vec{k}\shortminus\vec{q})}\dot{A}^{\sigma}_{\vec{q}}\Big{(}(\vec{k}-\vec{q}\,)\times\vec{\varepsilon}^{\,\,\sigma}_{\vec{q}}\Big{)},$
@@
$\big{(}\pi_{\phi}\vec{B}\,\big{)}_{\vec{k}}^{-}\equiv-\,i\int d^{3}\vec{q}\,\,\dot{\phi}^{*}_{(\vec{k}\shortminus\vec{q})}\,q\,\Big{(}A^{+}_{\vec{q}}(\vec{\varepsilon}^{\,+}_{\vec{k}}\cdot\vec{\varepsilon}^{\,+}_{\vec{q}})\shortminus A^{-}_{\vec{q}}(\vec{\varepsilon}^{\,+}_{\vec{k}}\cdot\vec{\varepsilon}^{\,-}_{\vec{q}})\Big{)}$
[…]
where we have used $\vec{\varepsilon}^{\,\,+}_{\vec{k}}\cdot\vec{\varepsilon}^{\,\lambda}_{\vec{q}}=(1-\lambda\cos\theta)/2$, and in the second line we have exploited the fact that $|A_{\vec{q}}^{+}|\gg|A_{\vec{q}}^{-}|$ holds during the initial stages of non-linear dynamics, over the region of excitation. Similarly, we can obtain analogous expressions for $\big{(}\pi_{\phi}\vec{B}\,\big{)}_{\vec{k}}^{\parallel}\,\equiv\,\hat{k}\cdot\big{(}\pi_{\phi}\vec{B}\,\big{)}_{\vec{k}}$, which will source $A_{\vec{k}}^{\parallel}$, and also for $\big{(}\vec{\nabla}\phi\times\vec{E}\,\big{)}_{\vec{k}}^{-}\,\equiv\,\vec{\varepsilon}^{\,\,+}_{\vec{k}}\cdot\big{(}\vec{\nabla}\phi\times\vec{E}\,\big{)}_{\vec{k}}$ and $\big{(}\vec{\nabla}\phi\times\vec{E}\,\big{)}_{\vec{k}}^{\parallel}\,\equiv\,\hat{k}\cdot\big{(}\vec{\nabla}\phi\times\vec{E}\,\big{)}_{\vec{k}}$, which will further source $A_{\vec{k}}^{-}$ and $A_{\vec{k}}^{\parallel}$, respectively. Remarkably, this means that the negative and longitudinal components of the gauge field will be sourced to high large amplitudes by the exponentially excited tachyonic modes $A_{\vec{q}}^{+}$, as e.g. shown explicitly by the expression in the second line of Eq. (V.3).
@@
Conversely, for a homogeneous inflaton, $\vec{q}$ collapses into $\vec{k}$, and there is no power transfer between different components. In Eq. (V.3) this can be understood manisfestly as $\cos\theta\rightarrow 1$ in the second line, whereas more generically, all components $\big{(}\pi_{\phi}\vec{B}\,\big{)}_{\vec{k}}^{-}$, $\big{(}\pi_{\phi}\vec{B}\,\big{)}_{\vec{k}}^{\parallel}$, $\big{(}\vec{\nabla}\phi\times\vec{E}\,\big{)}_{\vec{k}}^{-}$, and $\big{(}\vec{\nabla}\phi\times\vec{E}\,\big{)}_{\vec{k}}^{\parallel}$ collapse to zero if the inflaton is homogeneous, because it holds that $\vec{\varepsilon}^{\,\,+}_{\vec{k}}\cdot\vec{\varepsilon}^{\,+}_{\vec{k}}=0$ and $\hat{k}\cdot\vec{\varepsilon}^{\,\pm}_{\vec{k}}=0$. In other words, only if the inflaton presents inhomogeneities, will the inflaton-gauge interactions terms act as sources for the longitudinal and negative helicity components.
@@
We provide a quantitative description of the growth of the non-tachyonic modes in Fig. 17. There we plot the evolution of the integrated fractional contribution of the power spectrum corresponding to $A^{+}(\mathcal{N},k)$, $A^{-}(\mathcal{N},k)$ and $A^{\parallel}(\mathcal{N},k)$, for $\alpha_{\Lambda}=14$ (top) and $\alpha_{\Lambda}=18$ (bottom). The total contribution is largely dominated by the (+)(+)( + ) component, but there is an interesting interplay between the ()(-)( - ) and the longitudinal component. The former remains near the vacuum solution until the inflaton inhomegeneities begin to affect its evolution, at around $\mathcal{N}\sim-0.5$ (top) and $\mathcal{N}\sim-1.5$ (bottom), and then starts growing exponentially, reaching rapidly amplitudes way above the vacuum solution. At those moments the longitudinal mode, which was vanishing during the linear regime, grows also exponentially, reaches the level of the negative helicity, and keeps growing at the same rate. Both $A^{-}(\mathcal{N},k)$ and $A^{\parallel}(\mathcal{N},k)$ are driven by the tachyonic growth of $A^{+}(\mathcal{N},k)$, once the inflaton inhomogeneities become noticeable, soon after the onset of non-linearities in the dynamics. The growth of the relative contributions stabilize at around $15-20\%$ (longitudinal component) and $2-3\%$ (negative helicity) during the (electro)magnetic slow-roll regime, and persists like that till the end of inflation (vertical line), for both couplings.

### v2

@@ IV.1 Initial condition on the lattice
In order to run any simulation, we need first to initialize all the fields in the lattice. In the case of the gauge field, we know that the Bunch-Davies (BD) vacuum (23) is the solution that sets the initial condition for each helicity component, for wavelengths well inside the Hubble radius. Such condition characterizes an initial quantum vacuum state $|0\rangle$ (for which $\hat{a}_{\bf k}|0\rangle=0$), whereas solving for Eqs. (31)-(33) [or more precisely for their lattice counterparts Eqs. (44)-(45) and (47)] assumes implicitly the use of classical fields. Given the quantum expectation value $s_{A}^{2}(\tau,{\mathbf{k}})\equiv\langle 0|\hat{A}^{\lambda}(\tau,{\bf k})\hat{A}^{\lambda^{\dagger}}\hskip-1.99168pt(\tau,{\bf k})|0\rangle\equiv|\mathcal{A}^{\lambda}(\tau,{\bf k})|^{2}$ for the chirality $\lambda$, the BD condition implies $s_{A}^{2}\longrightarrow{1\over 2k}$. This corresponds to initially vanishing occupation numbers, $n_{\bf k}=0$, however, we know that one of the gauge field chiralities ($\lambda=+$ in our convention) will grow exponentially fast during the linear regime. This leads, in turn, to exponentially growing occupation numbers, $n_{\bf k}\gg 1$, and as a result, we can actually approximate the vacuum expectation values of products of field operators, by ensemble averages of random fields, see e.g. [98]. This corresponds to the classical limit of the excited field(s).
@@
[…]
@@ IV.2 Linear regime on the lattice
$\bullet$ An educated guess of the window $k_{\rm IR},k_{\rm UV}]$ needed to capture at least the modes for which the gauge field power spectrum is significantly above the BD tail, assuming linear evolution of such modes until the end of slow-roll inflation.
@@
It can be observed that initially there is a phase where the BD tail dominates over the tachyonic excitation in $\rho_{\rm EM}$. While $\rho_{\text{EM}}$ decreases with the expansion, IR modes also get excited, and eventually the excited IR range of the spectrum comes to dominate over the UV vacuum tail. In this example, this happens around $\mathcal{N}\sim-3$. Before that, the relative difference between methods is of the order of $\sim 0.1\%$, which is caused by the randomness of the BD condition in the lattice. We have checked that increasing the accuracy of the integrator does not reduce this relative error. Once the excited part of the gauge spectrum dominates over the vacuum tail, the relative difference increases, reaching approximately $\sim 10\,\%$ depending on the random realization. This is due to the cosmic variance problem already noted for the difference between spectra in the IR modes in Fig. 1. We highlight this issue by including the $1\sigma$ deviation in shaded, and we show that the variability of the fluctuation in the IR modes are causing this effect. Extending the evolution to $\mathcal{N}=0$ (while assuming inflaton slow-roll dominance), the mean relative error decreases again down to $\sim 1\,\%$, as the peak shifts to intermediate scales in the range $k_{\text{IR}},k_{\text{UV}}]$, and the weight of the most discrepant IR bins becomes less significant. In summary, regardless of the seed for the BD fluctuations, we correctly capture the evolution of the electromagnetic energy density in the lattice.
@@ IV.3 Switching to non-linear dynamics
For the largest couplings considered, in order to prevent the unwanted contribution of the gauge field vacuum tail to affect the dynamics, an additional step is required. Due to the special sensitivity to the UV of the non-linear regime, the lattice maximum momentum $k_{\rm UV}$ (or the comoving spatial resolution) needs to be sufficient to cover all relevant modes that will be excited in such regime. If the BD vacuum solution was imposed to all those modes, the non-classical contribution to the electromagnetic backreaction terms would be too large and would become the evolution unphysical. We tackle this issue introducing an intermediate cutoff $k_{\text{IR}}<k_{\text{BD}}<k_{\text{UV}}$, so that only within the ‘IR range’ $k_{\rm IR},k_{\rm BD}]$ we set the BD initial condition for the modes that are expected to be excited during the linear regime. Within the ‘UV range’ $k_{\rm BD},k_{\rm UV}]$, instead, the gauge field and electric field modes are set to zero during the linear dynamics, as they are not expected to become excited from the tachyonic instability. The gauge field modes within $k_{\rm BD},k_{\rm UV}]$ grow only stimulated out of the non-linear dynamics, when $\mathcal{N}>\mathcal{N}_{\rm switch}$. As a side effect of this procedure we can also tolerate to choose a larger time step $d\mathcal{N}$. We have performed numerical checks in order to ensure that applying such a cutoff does not alter the subsequent non-linear evolution, see App. D for a detailed discussion.
@@
The different purple lines correspond to different integration ranges for the integrated electromagnetic contributions. The dark solid lines correspond to the contribution calculated in the range $k_{\text{IR}},k_{\text{BD}}]$ and the light ones in $k_{\text{IR}},k_{\text{UV}}]$. While the former is the contribution we measure in our simulations, the latter is what we would have if we set the BD solution for the whole dynamical range. The non-continuous lines correspond to the isolated contribution of the IR excitation, that is above the BD initial condition of the gauge spectra. We do this by integrating the spectra up to a time-increasing cutoff $k_{\rm cut}(\mathcal{N})$ that grows from $k_{\rm IR}$ up to $k_{\rm BD}$. The dashed lines correspond to the range $k_{\rm IR},k_{\rm cut}(\mathcal{N})]$, whereas the dotted lines to $k_{\rm min},k_{\rm cut}(\mathcal{N})]$, where $k_{\text{min}}$ corresponds to Hubble crossing of the CMB scales at $\mathcal{N}=-60$, and with which we capture using the 1D method the gauge excitation during the whole cosmic history.
@@
Additionally, we see that the IR excitation captured by our lattice momentum window (dashed lines) is initially very subdominant, but it eventually becomes comparable to the vacuum tail’s contribution around $\mathcal{N}=-1.5$. Furthermore, it is also necessary that by the time we reach $\mathcal{N}_{\text{switch}}$, the IR excitation on the lattice dominates over that corresponding to modes below our IR threshold, i.e. $k<k_{\text{IR}}$. We assess this by including the contribution of the excited part computed in the range $k_{\rm min},k_{\rm cut}(\mathcal{N})]$ (dotted lines). The comparison with the dashed lines demonstrates that by the time the tachyonic excitation dominates over the UV tail, it also acquires well the level of excitation that considers the full cosmic history.
Figure 3: Comparison for $\alpha_{\Lambda}=15$ (left) and $18$ (right) of the terms related to the backreaction used to establish the time $\mathcal{N}_{\text{switch}}$. Top: the evolution of $\rho_{\rm K}$ (red), $\rho_{\rm V}$ (black) and $\rho_{\rm EM}$ (purple). Bottom: the evolution of $|\dot{\pi}_{\phi}|$ (red), $3H|\pi_{\phi}|$ (orange), and $m^{2}|\phi|$ (black) of Eq. (5), together with $\frac{\alpha_{\Lambda}}{a^{3}m_{p}}|\langle\vec{E}\cdot\vec{B}\rangle|$ (purple). The different purple lines represent different integration ranges: the dark solid purple lines are the contribution integrated for the range $k_{\text{IR}},k_{\text{BD}}]$, the light purple the integral for $k_{\text{IR}},k_{\text{UV}}]$, the dashed line is for only the excited IR contribution for $k_{\rm IR},k_{\rm cut}(\mathcal{N})]$, and the dotted purple line is the contribution of the excitation considering the whole cosmic history, i.e. $k_{\rm min},k_{\rm cut}(\mathcal{N})]$ (see text). The vertical dashed gray lines correspond to the times $\mathcal{N}_{\text{switch}}$.
@@
Finally, we note that the effect of the action of the cutoff scale $k_{\rm BD}$ is more pronounced for the largest couplings. As in the previous example, we include the electromagnetic components computed considering that we set the BD vacuum solution all across the range $k_{\text{IR}},k_{\text{UV}}]$ using light solid purple lines. It clearly shows that without the cutoff at $k_{\rm BD}$ an incredible overestimation of the integrated contributions will occur, which would falsify completely the dynamics. We conclude that for the largest coupling the use of an intermediate cutoff is compulsory.
@@ IV.4 Homogeneous Backreaction
$\alpha_{\Lambda}$ N()N\ ($\geq$) $k_{\rm{IR}}/m$ $k_{\rm{UV}}/m\ (\geq)$ $k_{\rm{BD}}/m$ $\mathcal{N}_{\rm{start}}$ $\mathcal{N}_{\rm{switch}}$
@@
Table 1: Summary of the relevant parameters of our simulations, with $\alpha_{\Lambda}$ the dimensionless coupling constant, NN the number of lattice sites per dimension, $k_{\rm{IR}}$ the infrared momentum, $k_{\rm{UV}}$ the ultraviolet momentum, $k_{\rm{BD}}$ the intermediate cutoff scale, $\mathcal{N}_{\rm{start}}$ the efold at which we start the simulation in the linear regime and $\mathcal{N}_{\rm{switch}}$ the efold in which we evolve the full equations of motion. For each coupling, only the cases with the largest separation of scales (or NN) are included.
@@
${\rm Linear}:$ $\hskip-5.69054pt\Delta\mathcal{N}_{\rm{br}}=m_{2}(\alpha_{\Lambda}-14.31)+1.44\,$, (79)
$\hskip-14.22636pt{\rm Power-law}:$ $\hskip-5.69054pt\Delta\mathcal{N}_{\rm{br}}=b_{2}(\alpha_{\Lambda}-14.31)^{a_{2}}+1.44\,$, (80)
@@
Figure 9: The energy components normalised to the total energy density, for the couplings $\alpha_{\Lambda}=12,\ 14$ and $18$, representative, respectively, of the weak, mild and strong backreaction regimes. Colours represent different components: black for potential, red for kinetic, blue for gradients, and purple for electromagnetic. The solid vertical lines corresponds to the end of inflation in each case. The dash-dotted line for $\alpha_{\Lambda}=14$ corresponds to the re-entering into the inflationary period.
@@
In Fig. 9 we show the evolution of different energy contributions for couplings representative of each regime, $\alpha_{\Lambda}=12,\,14$ and $18$, which correspond (from left to right in the figure), to the weak, mild and strong coupling regimes, respectively. More specifically, we plot the evolution of different energy densities with respect the total one, where $\rho_{\rm{K}}/\rho_{\rm tot}$ is depicted in red, $\rho_{\rm{V}}/\rho_{\rm tot}$ in black, $\rho_{\rm{G}}/\rho_{\rm tot}$ in blue and $\rho_{\rm{EM}}/\rho_{\rm tot}$ in purple. The vertical grey solid lines in each panel indicate the point where $\epsilon_{H}=1$, signalling the end of inflation for each coupling.
@@ V.2 Inflating UV scales
We observe that although there is a convergence in $\epsilon_{H}$ in the initial phase of backreaction, $2\lesssim\mathcal{N}\lesssim 2$, trajectories start diverging during the final stages towards the end of inflation. Simulations with the worst resolution (smaller $k_{\rm UV}/k_{\rm IR}$ ratio) deviate earlier, whereas the ones with better resolution sustain convergence up to $\mathcal{N}\sim 3.5$, and then diverge afterwards.
@@ V.3 Partial restoration of chirality balance and loss of transversality
$\big(\pi_{\phi}\vec{B}\,\big)_{\vec{k}}=-i\sum_{\lambda=\pm}\lambda\int\hskip-1.42262ptd^{3}q\,\,\dot{\phi}^{*}_{(\vec{k}\shortminus\vec{q})}\,q\,A^{\lambda}_{\vec{q}}\,\vec{\varepsilon}^{\,\lambda}_{\vec{q}}\,$, (84)
$\hskip-14.22636pt\big(\vec{\nabla}\phi\times\vec{E}\,\big)_{\vec{k}}=i\sum_{\sigma=\pm\,,\parallel}\int\hskip-1.42262ptd^{3}q\,\,\phi^{*}_{(\vec{k}\shortminus\vec{q})}\dot{A}^{\sigma}_{\vec{q}}\Big((\vec{k}-\vec{q}\,)\times\vec{\varepsilon}^{\,\,\sigma}_{\vec{q}}\Big),$
@@
$\hskip-14.22636pt\big(\pi_{\phi}\vec{B}\,\big)_{\vec{k}}^{-}\equiv-\,i\int\hskip-1.42262ptd^{3}\vec{q}\,\,\dot{\phi}^{*}_{(\vec{k}\shortminus\vec{q})}\,q\,\Big(A^{+}_{\vec{q}}(\vec{\varepsilon}^{\,+}_{\vec{k}}\hskip-2.84526pt\cdot\hskip-1.42262pt\vec{\varepsilon}^{\,+}_{\vec{q}})\shortminus A^{-}_{\vec{q}}(\vec{\varepsilon}^{\,+}_{\vec{k}}\hskip-2.84526pt\cdot\hskip-1.42262pt\vec{\varepsilon}^{\,-}_{\vec{q}})\Big)$
$\hskip 25.03856pt\simeq-\,{i\over 2}\int\hskip-0.85358ptd^{2}\hat{q}\,dq\,q^{3}\,\,\dot{\phi}^{*}_{(\vec{k}\shortminus\vec{q})}\,(1-\cos\theta)\,A^{+}_{\vec{q}}\,$, (86)
where we have used $\vec{\varepsilon}^{\,\,+}_{\vec{k}}\hskip-1.42262pt\cdot\vec{\varepsilon}^{\,\lambda}_{\vec{q}}=(1-\lambda\cos\theta)/2$, and in the second line we have exploited the fact that $|A_{\vec{q}}^{+}|\gg|A_{\vec{q}}^{-}|$ holds during the initial stages of non-linear dynamics, over the region of excitation. Similarly, we can obtain analogous expressions for $\big(\pi_{\phi}\vec{B}\,\big)_{\vec{k}}^{\parallel}\,\equiv\,\hat{k}\cdot\big(\pi_{\phi}\vec{B}\,\big)_{\vec{k}}$, which will source $A_{\vec{k}}^{\parallel}$, and also for $\big(\vec{\nabla}\phi\times\vec{E}\,\big)_{\vec{k}}^{-}\,\equiv\,\vec{\varepsilon}^{\,\,+}_{\vec{k}}\cdot\big(\vec{\nabla}\phi\times\vec{E}\,\big)_{\vec{k}}$ and $\big(\vec{\nabla}\phi\times\vec{E}\,\big)_{\vec{k}}^{\parallel}\,\equiv\,\hat{k}\cdot\big(\vec{\nabla}\phi\times\vec{E}\,\big)_{\vec{k}}$, which will further source $A_{\vec{k}}^{-}$ and $A_{\vec{k}}^{\parallel}$, respectively. Remarkably, this means that the negative and longitudinal components of the gauge field will be sourced to high large amplitudes by the exponentially excited tachyonic modes $A_{\vec{q}}^{+}$, as e.g. shown explicitly by the expression in the second line of Eq. (V.3).
@@
Conversely, for a homogeneous inflaton, $\vec{q}$ collapses into $\vec{k}$, and there is no power transfer between different components. In Eq. (V.3) this can be understood manisfestly as $\cos\theta\rightarrow 1$ in the second line, whereas more generically, all components $\big(\pi_{\phi}\vec{B}\,\big)_{\vec{k}}^{-}$, $\big(\pi_{\phi}\vec{B}\,\big)_{\vec{k}}^{\parallel}$, $\big(\vec{\nabla}\phi\times\vec{E}\,\big)_{\vec{k}}^{-}$, and $\big(\vec{\nabla}\phi\times\vec{E}\,\big)_{\vec{k}}^{\parallel}$ collapse to zero if the inflaton is homogeneous, because it holds that $\vec{\varepsilon}^{\,\,+}_{\vec{k}}\hskip-1.42262pt\cdot\vec{\varepsilon}^{\,+}_{\vec{k}}=0$ and $\hat{k}\hskip-1.42262pt\cdot\vec{\varepsilon}^{\,\pm}_{\vec{k}}=0$. In other words, only if the inflaton presents inhomogeneities, will the inflaton-gauge interactions terms act as sources for the longitudinal and negative helicity components.
@@
We provide a quantitative description of the growth of the non-tachyonic modes in Fig. 17. There we plot the evolution of the integrated fractional contribution of the power spectrum corresponding to $A^{+}(\mathcal{N},k)$, $A^{-}(\mathcal{N},k)$ and $A^{\parallel}(\mathcal{N},k)$, for $\alpha_{\Lambda}=14$ (top) and $\alpha_{\Lambda}=18$ (bottom). The total contribution is largely dominated by the $(+)$ component, but there is an interesting interplay between the ()(-) and the longitudinal component. The former remains near the vacuum solution until the inflaton inhomegeneities begin to affect its evolution, at around $\mathcal{N}\sim-0.5$ (top) and $\mathcal{N}\sim-1.5$ (bottom), and then starts growing exponentially, reaching rapidly amplitudes way above the vacuum solution. At those moments the longitudinal mode, which was vanishing during the linear regime, grows also exponentially, reaches the level of the negative helicity, and keeps growing at the same rate. Both $A^{-}(\mathcal{N},k)$ and $A^{\parallel}(\mathcal{N},k)$ are driven by the tachyonic growth of $A^{+}(\mathcal{N},k)$, once the inflaton inhomogeneities become noticeable, soon after the onset of non-linearities in the dynamics. The growth of the relative contributions stabilize at around $15-20\%$ (longitudinal component) and $2-3\%$ (negative helicity) during the (electro)magnetic slow-roll regime, and persists like that till the end of inflation (vertical line), for both couplings.